I need help on some homework please and thank you.
Given the following measurements, A=42degrees, a=44cm, and c=58cm how many solutions are possible?

Answer by stanbon(75887)

I need help on some homework please and thank you.
Given the following measurements, A=42degrees, a=44cm, and c=58cm how many solutions are possible?
-----
sin(C)/58 = sin(42)/44
---
sin(C) = (58/44)sin(42) = .8820
C = 61.89 degrees OR C= 180-61.89 = 118.12 degrees
----
Then B = 180-(42+61.89) = 76.11 degrees
OR
B = 180-(42+118.12) = 19.88 degrees
======================================
Cheers,
Stan H.
============